OpenStudy (anonymous):

Which of the following is the equation of a circle whose center is at the origin and whose radius is square root of 5 ?

OpenStudy (anonymous):

x2 + y2 = 5 x2 + y2 = 10 x2 + y2 = 25

OpenStudy (chriss):

The equation of a circle with the center at the origin is\[x^{2}+y^{2}=r^{2}\]